OpenStudy (anonymous):

solve the equation by factoring x^2+7x-44=0

OpenStudy (anonymous):

x^2+7x-44=0 Or, x^2+11x-4x-44=0 Or, x^2-4x+11x-44=0 Or, x(x-4)+11(x-4)=0 Or, (x-4)(x+11)=0 Therefore, x=4, -11

OpenStudy (lalaly):

(x-4)(x+11)